eric averaged 50 MPH traveling from denevr to minneapolis. Returning by a different route that covered the same number of miles, he averaged 55 mph. what is his distance between the two cities if hid total travel time was 32 hrs?
going: r=50 t=x
returning: r=55 t=32-x
i dont know how to even begin.

Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

This is the rate*time=distance problem
we are given
t1 + t2 = 32
also
50t1 = d and
55t2 = d
we have 50t1 = 55t2 and
t1 = 55t2 / 50
substitute for t1 in the first equation and we get
55t2/50 + t2 = 32
multiply both sides of the = by 50
55t2 + 50t2 = 1600
105t2 = 1600 and t2 = 15.2
distance = 55 * 15.2 = 836 miles

Answer by Alan3354(69443)

eric averaged 50 MPH traveling from denevr to minneapolis. Returning by a different route that covered the same number of miles, he averaged 55 mph. what is his distance between the two cities if hid total travel time was 32 hrs?
------------
Avg speed for the round trip = 2*50*55/(50+55) = 5500/105
= 1100/21 mi/hr
-----------
RT distance = (1100/21)*32
One way = 17600/21 miles
=~ 838.1 miles